Q: Evaluate the following integral using residue theorem

∫z^2 e^z dz ; C : |z| = 1


1
Expert's answer
2020-10-28T18:30:03-0400

Consider the integral "\int_C z^2e^z" where "C=\{z\in\mathbb C: |z|=1\}". The function "f(z)=z^2e^z" has no a singular point, and is a holomorphic function in "D=\{z\in\mathbb C: |z|\leq 1\}". Therefore, Cauchy’s residue theorem implies "\int_C z^2e^z =0."